Find [Cu2+] in a solution saturated with Cu4(OH)6(SO4) if [OH-] is fixed at 6.5 x 10-6 M. Note that Cu4(OH)6(SO4) gives 1 mol of SO42- for 4 mol Cu2+. Cu4(OH)6(SO4)(s) <--> 4Cu2+ + 6OH- + SO42- Ksp = 2.3 x 10-69
Answer #1

Let the concentration of Cu2+ be x M. we know that 1 mol of SO42- give 4 mol of  Cu2+ . So concentration of SO42- = x/4

We know, for a rxn: aA   bB + cC

Ksp = ( [B]b [C]c ) / [A]a

Solids are not included when calculating equilibrium constant expressions, because their concentrations do not change the expression; any change in their concentrations are insignificant, and therefore omitted.

Ksp = [x4 X (x/4) X (6.5X10-6) 6] / 1 = 2.3 X 10-69

x5 = (2.3 X 10-69 ) / (18854.7226 X 10-36 ) = 1219.85 X 10-40

x = 4.1425 X 10-8 M = [Cu2+]

hope this helps! kindly upvote :)